OpenStudy (anonymous):

10. If tan(x) =4/3 and x is in quadrant III, find cos(x).

OpenStudy (dumbcow):

set up a triangle in quad3 use SOH CAH TOA tan = opp/adj  From the triangle, find cos cos = adj/hyp --> cos = -3/5